OpenStudy (anonymous):

what is -5/18 + 1/6? a. -4/9 b. -1/9 c. 1/9 d. 4/9

OpenStudy (lilg132):

b

OpenStudy (espex):

When working with fractions they must have a common denominator if you wish to add/subtract them to/from each other. Notice here that 6 is a factor of 18, so you can give them both a common denominator of 18 to do your math:\[\frac{-5}{18}+(\frac{1}{6} \times \frac{3}{3}) \rightarrow \frac{-5}{18}+\frac{3}{18}\] You can do this because if you multiply a number by 1 it doesn't change the number. Now you are able to carry out the arithmetic function on the numerators, -5+3, or 3-5 which is a -2. So your fraction now looks like this:\[\frac{-2}{18}\] Check to see if it is in its simplest form, which it is not, so you reduce. Multiply the top and bottom by 2, cancel, and you get \[\frac{-2}{18} \times \frac{2}{2}=\frac{-1}{9}\] So lilg32 was correct, b is your answer.